OpenStudy (anonymous):

can somebody help with -(6-7b)-6=2

OpenStudy (anonymous):

You want to isolate b. Factor in the negative and you get: \[-6+7b -6 = 2\]\[7b = 2+6+6\]\[7b=14\]\[b=2\]

OpenStudy (anonymous):

-(6-7b)-6=2 First distribute the negative sign -6+7b-6 =2 then isolate b 7b = 6+6+2 7b = 14 now solve for b b = 2

OpenStudy (anonymous):

thank you